OpenStudy (rianiscool):

need help asap will give a medal Jorge's rectangular garden was 17 feet long and had a perimeter of 54 feet. He made his garden 5 feet longer. What is the area of his new garden? A. 22 square feet B. 64 square feet C. 85 square feet D. 220 square feet

OpenStudy (anonymous):

Area = l * w, Perimeter = l+l+w+w or 2l+2w. Since you know the garden is a rectangle then you know two of the side lengths are 17. If you plug this into the perimeter formula, along with the given perimeter you get 54=2(17)+2(w). Then if you solve for w you will get the other side lengths of the rectangle. You can use that length and the 17+5 to find area of the new rectangle.

OpenStudy (anonymous):

So 54=2(17)+2(w) 54=34+2(w) 54-34=2(w) 20=2(w) divide both sides by 2 to get 10=w. Lets check to see if this is right: Does 54= 2(17)+2(10)? yes it does! So we know the length of the first rectangle is 17 and the width is 10. Then we have to add on 5 ft to the length so 17+5=22. 22 is the length of the new rectangle/garden. We still have 10 as the width. So to find the area of the new garden we just use the area formula A=l*w A=22*10 Area=220 SO the answer is D. 220 square feet

OpenStudy (rianiscool):

@StudyGurl14

OpenStudy (rianiscool):

@iGreen need help is it d

OpenStudy (studygurl14):

I believe so, yes. @Runner123456789 seems to have done a really good job showing you the steps. :)